the product of two consecutive integers is 3 less than 3 times their sum. Find the integers.: the product of two consecutive integers is 3 less than 3 times their sum. Find the integers.
Answer by scott8148(2759)

"the product of two consecutive integers" __ x(x+1) __ x^2+x

"is" __ =

"3 less than 3 times their sum" __ 3(2x+1)-3 __ 6x

x^2+x=6x __ subtracting 6x __ x^2-5x=0 __ factoring __ x(x-5)=0

x=0 __ 0 and 1

x-5=0 __ x=5 __ 5 and 6